In a colony all families have at least two child.The probability that a randomly chosen family from this colony has exactly k children is
4/5*(1/5)^k-2;k=2,3... A child is either a male or a female with equal
probability.What will be the probability that a randomly selected family has
EXACTLY 2 boys?

Answer by Edwin McCravy(20060)

In a colony all families have at least two child.The probability that a randomly chosen family from this colony has exactly k children is
4/5*(1/5)^(k-2);k=2,3... A child is either a male or a female with equal
probability.What will be the probability that a randomly selected family has
EXACTLY 2 boys?
The probability of having EXACTLY 2 children is found by substituting 2 for k
in the formula

4/5*(1/5)^(2-2) = 4/5*(1/5)^0 = 45*1 = 4/5

There are 4 ways to have EXACTLY 2 children, BB, BG, GB, and GG.  So we must divide 4/5 by 4.

Answer: 1/5

[Note: Maybe you wonder why we didn't divide by 2 instead of 4.  That's
because there are just as many of one of the 4 cases as any of the others.]

Edwin